OpenStudy (chris215):

The base of a solid in the xy-plane is the circle x^2 + y^2 = 16. Cross sections of the solid perpendicular to the y-axis are semicircles. What is the volume, in cubic units, of the solid?

OpenStudy (chris215):

I got 2pi/3

OpenStudy (holsteremission):

This solid is exactly a hemisphere of radius \(4\). \[V=\frac{1}{2}\left(\frac{4}{3}\pi (4^2)\right)=\frac{32\pi}{3}\]

OpenStudy (chris215):

ohh ok tthtanks :)

OpenStudy (holsteremission):

Whoops, I meant \(\color{red}{(4^3)}\), should be \(\dfrac{128\pi}{3}\).

OpenStudy (holsteremission):

Are you supposed to find the volume using integration?

OpenStudy (chris215):

Im not sure

OpenStudy (chris215):

I think the answer is 32π/3

OpenStudy (holsteremission):

Why do you say that?

OpenStudy (chris215):

never mind tht does seem right

OpenStudy (holsteremission):

Right. In case you need more convincing (and judging by the fact that you seem to have posted several calc question recently), here's a justification using integration.  Each section's radius is given by the vertical distance to the horizontal axis, which is given by the positive root \(\sqrt{16-x^2}\). The area of any one cross section is \(\dfrac{\pi r(x)^2}{2}\), with \(r(x)\) the radius which depends on the section's position along the axis. The total volume is the sum of infinitely many, infinitesimally thin sections, so you have \[V=\int_{-4}^4 \frac{\pi(\sqrt{16-x^2})^2}{2}\,\mathrm{d}x\]which you'll find to coincide with \(\dfrac{128\pi}{3}\).